### v2

# SLAC-PUB-251124 January 2026

ABSTRACT
In this lecture, I discuss measurements of the properties of the Higgs boson and related observables in the era of Higgs factories. This highly motivated experimental program is the challenge for the next generation of particle physicists.
Invited plenary lecture presented at the
32nd International Symposium on Lepton Photon Interactions at High Energies,
Madison, Wisconsin, USA, August 25-29, 2025
1 Introduction
The Higgs boson has always been at the center of the Standard Model of Particle Physics. Without it, the physics of the Standard Model would be impoverished, with no spectroscopy or flavor physics, the universe a dilute soup of massless mesons and leptons. The discovery of the Higgs boson at the Large Hadron Collider in 2012[1, 2] gave a crucial verification of this important idea. Since this discovery, the LHC experiments ATLAS and CMS have demonstrated, for the $W$ and $Z$ bosons and for an increasing number of the quarks and leptons, that the Higgs field provides at least the major source of their masses [3, 4].
These experimental advances ought to change our perspective on the study of particle physics. We now see clearly that the major outstanding questions about particle physics are questions about the interactions of the Higgs boson. We need to learn more about this particle, and we have the resources to gain that knowledge. We will learn much from the High Luminosity phase of the LHC, but, beyond this, we will be able to measure the properties of the Higgs boson with high precision from a future “Higgs Factory”, an $e^{+}e^{-}$ collilder operating at energies from the $Z$ boson resonance to the top quark threshold and above. This has been recognized in the latest strategy statements from all of the major world regions in particle physics. From the 2020 Update of the European Strategy for Particle Physics: “An electron-positron Higgs factor is the highest priority next collider.” [5]. From the 2022 report of the Japan Association of High Energy Physicists (JAHEP): “An $e^{+}e^{-}$ Higgs Factory is the most important next large-scale particle physics facility”[6]. From the 2023 US High Energy Physics Advisory Panel P5 report: “The next step is to use electron and positron beams to construct a Higgs factory, which would allow precision measurements of the Higgs boson properties and searches for exotic decays, possibly into dark matter” [8]. In China, the Circular Electron-Positron Collilder (CEPC) was named the highest priority next project in particle and nuclear physics in the first round of CAS deliberations [9], though eventually the CAS chose to endorse a less ambitious collider. Very recently, the European Strategy Group has recommended the FCC-ee, a circular Higgs factory, as the next flagship CERN project [10].
In this lecture, I review the importance of the measurements expected from Higgs factories and the levels of precision that they might achieve. Section 2 reviews the central role of the Higgs boson that motivates these measurements. In Sections 3–5, I discuss Higgs measurements in three different programs avaiable at Higgs factories. Section 3 discusses the program at the peak of the $e^{+}e^{-}\to HZ$ cross section at 240-250 GeV in the $e^{+}e^{-}$ center of mass. Section 4 discusses the program of precision measurements at the $Z$ resonance. Section 5 discusses measurements above the top quark threshold. All three sections touch on the question of the comparison of linear and circular $e^{+}e^{-}$ colliders. The European Strategy Group has now recommended the FCC-ee as the next CERN project, but still it is interesting to summarize what we have learned about this comparison in the past year of physics discussions. Section 6 discusses the fascinating technical opportunities that the Higgs factory physics program will provide. Section 7 will give some conclusions.
The discussion in Sections 3–5 relies heavily on the excellent work reported in the Physics Briefing Book for the 2026 update of the European Strategy for Particle Physics [11] and, in particular, on the new global fit defining the capabilities of future colliders. This is a major update of the global analysis prepared for the Snowmass 2022 study [12].
The discussions for the update of the European Strategy for Particle Physics have considered both the goal of improving our knowledge of the Higgs boson and the goal of providing a collider that can reach 10 TeV in the parton center of mass. In this paper, I will concentrate on the first of these aims. This is the program that needs to be defined now — the physics of precision Higgs boson measurements available with current accelerator technologies and the strategies for achieving the goals of this program. As I will emphasize in this paper, this is the most important goal concerning future colliders for the progress of our field and our commitment to the current generation of young physicists. In parallel, we must study the choices for accelerators reaching 10 TeV parton collisions. But today there is no affordable technology that meets this goal, and it will take decades of accelerator R&D to develop and evaluate possible paths.
2 Importance of studying the Higgs boson with high precision
Yes, what I said above is correct. The major outstanding questions about particle physics are in fact questions about the properties of the Higgs boson.
The Standard Model (henceforth, SM) has two distinct pieces. The first is the gauge Lagrangian
${\cal L}_{gauge}=\sum_{a=1,2,3}-{1\over 4}(F^{a}_{\mu\nu})^{2}+\sum_{f}\overline{f}\ i\not{\hbox{\kern-4.0pt$D$}}\ f\ ,$ (1)
which describes the motion of gauge bosons, quarks, and leptons and their couplings to one another through a beautiful geometrical structure. This Lagrangian has only three parameters, the gauge couplings of $U(1)$, $SU(2)$, and $SU(3)$, whose values tempt us to believe that they derive from a single grand-unified coupling. The second is the Higgs Lagrangian,
${\cal L}_{Higgs}=|D_{\mu}\Phi|^{2}-V(\Phi)-\sum_{f_{L},f_{R}}Y_{f}\ \overline{f}_{L}\cdot\Phi f_{R}+(h.c.),$ (2)
where $V(\Phi)$ is the Higgs field potential and the third term is a shorthand for the quark and lepton Yukawa couplings.
In a renormalizable theory, $V(\Phi)$ has 2 parameters, the Higgs mass and self-coupling; if the SM is an effective theory, it can have many more. The Yukawa couplings have a total of 54 parameters, of which only 16 are observable if the SM is exact. Neutrino masses are easily accommodated by adding right-handed neutrinos $\nu_{R}$ with their own mass matrix and Yukawa interactions. The SM gives us no physical explanation for any of these couplings or any way of computing them from deeper principles. Yet, they contain a huge amount of physics – the spectrum of quark and lepton masses, the quark flavor mixings and CP violation , and, in the model with $\nu_{R}$, the origin of neutrino masses. The SM does not explain the preponderance of matter over antimatter in the universe, but any explanation would require CP violation due to scalar bosons, either the Higgs bosons or their possible heavy partners. The SM does not contain cosmic dark matter, and there are many ad hoc models of dark matter than do not invoke the Higgs sector. But all dark matter models with other physical motivations, in particular, explanations involving supersymmetry or axions, invoke either the Higgs boson or its partners.
Above all of these questions, there is the question of why the Higgs field behaves in its special way: Why does the Higgs field acquire a vacuum expectation value and break the symmetry of the SM? There are many examples of spontaneous symmetry breaking in physics, from superconductivity and magnetism in condensed matter systems to chiral symmetry breaking in QCD. In each of these cases, there is a compelling physical explanation for the fact that the system forms an asymmetric ground state. The Higgs symmetry breaking is the one exception. In the SM, the instability of the symmetrical vacuum state is put in by hand. Theorists have proposed many models of Higgs symmetry breaking, including models with supersymmetry, models in which the Higgs field is a composite state, and models in which the Higgs field arises from the physics of extra space dimensions. But at this moment there is no experimental evidence for any of them. Importantly, each of these approaches gives a different set of explanations of all of the phenomena listed in the previous paragraph. Without the answer to this question, we cannot make progress against those major problems of particle physics.
To summarize this situation: “Higgs isn’t everything. It’s the only thing”***Apologies to Red Sanders [13]..
The mystery of the origin of electroweak symmetry breaking is tightly connected to another often-discussed issue in particle physics, the “hierarchy problem” or the “naturalness” problem. These days, with the failure, so far, to discover new particles at the LHC, it has become fashionable not to believe in naturalness. I am very much opposed to this attitude. For an extended discussion, please see my recent paper on this question [14]. As a point of emphasis, I remind you that the naturalness of the electroweak scale is the only reason that we expect to see the breakdown of the SM at the 10 TeV scale, as opposed to a much higher and probably inaccessible energy scale. Would you build a 10 TeV collider at great expense if there is no expectation of a discovery at that machine?
Figure 1: The correspondence between Higgs boson couplings and particle mass, according to the Higgs coupling measurements achieved by the CMS experiment, from [4]. The lower plot shows the residual uncertainties on a linear scale.
There is one more issue concerning the Higgs boson that needs a comment. This is the truly impressive job that the ATLAS and CMS experiments have done in measuring the couplings of the Higgs boson, using observed Higgs boson decays and also the production of $t\overline{t}H$ final states. The results as of the end of run 2 of the LHC are collected in Fig. 1, from CMS, and a similar figure from ATLAS, both published in Nature [3, 4]. This data demonstrates that the Higgs boson vacuum expectation value is the main source of mass for the particles shown. Many people interpret these beautiful results as proof that the observed Higgs boson precisely follows the Standard Model. To a great extent, this is true, but it is important to think more carefully.
The SM relation that, for each particle, the Yukawa coupling is proportional to the particle mass, can be corrected by the effects of new particles from a beyond-SM (BSM) theory. But we should expect that these corrections are small. The Yukawa couplings are determined, at $q^{2}=0$, from the measured values of the masses. These couplings, at $q^{2}=m_{H}^{2}$, can be measured in Higgs decays New particles of heavy masses $M$ can generate a difference between the SM prediction and the measured value of the couplings, but this will be proportional to
$m_{H}^{2}/M^{2}\ .$ (3)
The fact that the LHC has not already discovered these new particles makes it very likely that their masses are above 1 TeV. Then the effects of BSM particles on Higgs couplings are expected to be at the level of a few percent. The same conclusion follows from more formal arguments based on Effective Field Theory or quantum field theory “decoupling”. This means that, however impressive the LHC results are, we are not yet in the game of observing BSM effects in the Higgs couplings.
The expected results from the High-Luminosity LHC will bring us to uncertainties of 2-4% – enough to suggest an effect, but likely not enough to prove it.
There is a point here that is often ignored in discussions of future precision measurements: The purpose of precision is discovery, not just improving the error bars. It is possible to make discoveries with precision, but we must realize that the burden of proof is much higher than for the discovery of resonances. The community will be skeptical about a precision measurement that deviates from the SM, and with good reason, considering the recent history of SM anomalies. It will be necessary to bring arguments that can overcome this skepticism. This requires multiple cross-checks on the measurement and the associated interpretation. Systematic errors, especially those obtained by human estimation, should be subdominant. It is especially important that the same deviation should be seen in different settings with different sources of uncertainty. I will return to these criteria as I discuss the discovery opportunities from various stages of the Higgs precision program,
3 Higgs measurements at threshold
At the moment, there are a number of Higgs factory proposals under consideration around the world. These include circular $e^{+}e^{-}$ colliders, the Future Circular Collider (FCC-ee), at CERN [15, 16, 17, 18], and the Circular Electron-Positron Collider (CEPC) in China [19, 20]. They also include linear $e^{+}e^{-}$ colliders, such as the ILC in Japan [21], the Compact Linear Collider (CLIC) at CERN [22], and the relatively newly proposed Linear Collider Facility (LCF) at CERN [23]. All of these, except for CEPC, have been studied as candidate future machines for the European Strategy for Particle Physics. For the details of the accelerator designs and estimated costs and schedules, please follow the references; there is no room for a full discussion here.
Figure 2: Expected instantaneous luminosities as a function of CM energy for proposed circiular and linear Higgs factories, from  [24].
Some crucial difference between the linear and circular designs need to be emphasized: These designs have luminosities that depend in very different ways on the CM energy. For the circular designs, the luminosity is limited by synchrotron radiation, leading to a very high luminosity at the lowest energies which then falls as $E_{CM}^{-4}$ as the CM energy is increased. Linear colliders have luminosities limited by beam-beam effects, leading to a luminosity increasing as $E_{CM}^{1}$, as is typical for lower-energy synchrotrons. The expected luminosities for many of the projects cited above are shown in Fig. 2 [24]. The crossover is at about 350 GeV, close to the $t\overline{t}$ threshold. The run plans for the various projects reflect this dependence. In particular, the luminosity for circular machines is about 3 orders of magnitudes higher than for linear machines at the $Z$ pole, but falls too low for an interesting physics program at CM energies of 400 GeV and above. Linear colliders can support longitudinal beam polarization both for electrons and positrons, an advantage in studying the chiral couplings of the SM. We will see how these differences play out as we consider the various energy stages of $e^{+}e^{-}$ Higgs factory operation.
Figure 3: Cross sections for Higgs boson production in $e^{+}e^{-}$ annhilation reactions as a function of CM energy, showing also the dominant Feynman diagrams for the reactions $e^{+}e^{-}\to ZH$ and $e^{+}e^{-}\to\nu\overline{\nu}H$.
Figure 4: Determination of the Higgs boson mass at 250 GeV from measurement of the $Z$ recoil mass in $e^{+}e^{-}\to ZH$ with $Z\to\mu^{+}\mu^{-}$, from [25]. The plot shown corresponds to an integrated luminosity of 250 fb-1, to be compared to 3 ab-1 for the LCF program at 250 GeV.
With this context, let’s begin our discussion of the expectations for the measurements of Higgs boson couplings – in particular, the couplings accessible in Higgs boson decay. The dominant cross sections for Higgs boson production at $e^{+}e^{-}$ colliders are Higgstrahlung, $e^{+}e^{-}\to ZH$, and $WW$ fusion, $e^{+}e^{-}\to\nu\overline{\nu}H$. The cross sections for these reactions in the CM energy range of Higgs factories is shown in Fig. 3. The cross sections for these two reactions cross over at about 470 GeV.
In this section, I will discuss the study of the Higgs from the Higgstrahlung reaction, optimized at the peak of the cross section at 240-250 GeV. This working point has important advantages for the Higgs study. At this point, each Higgs boson is tagged by a recoiling $Z$ boson with a lab-frame energy of 110 GeV. The peak in the $Z$ energy distribution, shown as a recoil mass of the Higgs boson, is presented in Fig. 4. Almost every $Z$ boson in the peak is recoiling against a Higgs boson observed in some possible decay mode. The background under the peak comes mainly from $e^{+}e^{-}\to ZZ$ plus radiation; it is smooth and also precisely calculable. Using only the observed $Z$ boson, it is possible to measure the absolute Higgstrahlung cross section and, using the best-defined $Z$’s with decay to $\mu^{+}\mu^{-}$, to measure the Higgs mass to 10 MeV ($10^{-4}$ precision). Looking on the other side of tagged events, it is possible measure ratios of branching ratios, including branching ratios to invisible and exotic modes, in an almost completely model-independent way.
The one missing piece of information from this study is the Higgs total width. The predicted width in the SM is 4 MeV, too small to be extracted from the recoil line shape. If all exotic decays of the Higgs (including invisible and partially-invisible modes) can be observed, the total width can be inferred from the absolute value of the Higgstrahlung cross section and the constraint that the branching ratios sum to 1. If one lacks confidence that all “unclassified” Higgs decay modes can be identified, it is still possible to determine the total width from a fit to some theoretical model.
Often, this model is provided by the kappa framework, in which each SM decay coupling $g(HAA)$ is multiplied by a factor $\kappa_{A}$. However, this approach is model-dependent and also gives some counterintuitive results. In typical fits, the $HZZ$ coupling has a tiny uncertainty while the $HWW$ coupling has an enhanced uncertainty. This defies our intuition, shaped by precision electroweak measurements, that electroweak interactions are tightly constrained by custodial symmetry that relates the couplings of $W$ and $Z$. The asymmetry is caused by the assumption of the kappa framework that the $HZZ$ coupling has the same Lorentz structure as in the SM, while, in general, additional structures can appear.
In [26], we suggested a better approach, based on Standard Model Effective Field Theory (SMEFT) [27]. SMEFT will appear several times in our discussion, so let me give a brief description of it here. SMEFT writes the true Lagrangian of nature as an expansion about the SM Lagrangian. The fields are taken to be those of the observed particles of SM, with possible heavier particles integrated out. The interactions are the most general ones that obey the SM gauge symmetries. The series is ordered by operator dimension,
${\cal L}={\cal L}_{SM}+\sum_{i}{c_{i}\over\Lambda^{2}}{\cal O}_{i}^{(6)}+\sum_{j}{d_{j}\over\Lambda^{4}}{\cal O}_{j}^{(8)}+\cdots\ .$ (4)
Here ${\cal L}_{SM}$ is the SM Lagrangian, which is actually the most general Lagrangian satisfying these criteria with only operators of dimension 4 and below. The additional terms listed are the contributions from operators of dimension 6 and 8. Operators of odd dimension necessarily violate lepton or baryon number, so I have omitted them for the discussion of Higgs properties. The parameter $\Lambda$ is the mass scale of new, Beyond-SM, particles. If $\Lambda>1$ TeV, the dimension 6 operators will already be small corrections to the SM and higher-dimension operators will be further suppressed.
A problem in working with SMEFT is that the total number of operators appearing at each level is large. For the SM with 3 generations, already at dimension 6 there are 2499 operators. However, for the analysis of Higgs couplings, assuming that electroweak perturbative corrections to the SMEFT corrections can be ignored, we can restrict ourselves to operators that appear in Higgs factory processes at the tree level, of which there are only 18. Adding two parameters for the Higgs invisible branching ratio (measured) and the “unclassified” branching ratio (assumed unmeasured) gives a closed fit. Since (4) is the full Lagrangian of nature, we can add precision measurements from other processes, including $e^{+}e^{-}\to W^{+}W^{-}$ and precision electroweak observables. With this very relevant information, the fit tightly constrains the Higgs boson total width and the absolute values of all couplings. This procedure can make use of beam polarization, if it is available, to resolve the coefficients of chiral operators; the kappa fit, with its restrictive assumptions, ignores the effect of beam polarization. Unsurprisingly, the use of beam polarization turns out to give a significant advantage.
coupling LCF 250 LCF 550 FCC-$e$ CEPC
bbbb 0.72 0.38 0.42 0.38
cccc 1.45 0.87 0.78 0.64
gggg 1.31 0.70 0.68 0.58
$\tau\tau$ 0.83 0.53 0.48 0.42
WWWW 0.34 0.22 0.26 0.28
ZZZZ 0.34 0.22 0.26 0.27
$\gamma\gamma$ 1.02 0.94 0.94 0.91
$\mu\mu$ 3.87 3.53 3.33 3.01
$\Gamma_{tot}$ 1.39 0.85 0.82 0.76
BRBR(invis.) 0.36 0.31 0.25 0.26
BRBR(other) 1.53 1.11 0.78 0.67
Table 1: Expected 1 $\sigma$ precisions ($\Delta g/g$, in %) for Higgs boson couplings from the expected run plans of linear and circular Higgs factories, from the SMEFT fit described in the text. The last two lines are 95% CL upper limits on branching ratios. The uncertainties in the $HWW$ and $HZZ$ couplings improve to about 0.15% if all exotic decays of the Higgs boson can be observed and there is no “unclassified” category of decays. The first two columns of this table were presented in [24].
coupling LCF 250 LCF 550 FCC-$e$
bbbb 0.59 0.27 0.30
gggg 0.82 0.53 0.44
$\tau\tau$ 0.69 0.45 0.38
WWWW 0.42 0.13 0.21
ZZZZ 0.50 0.13 0.21
$\gamma\gamma$ 0.82 0.53 0.44
Table 2: Expected 1 $\sigma$ precisions ($\Delta g/g$, in %) for Higgs boson couplings from the expected run plans of linear and circular Higgs factories, from the SMEFT fit presented in Figure 3.6 of [11]. This fit includes some NLO SMEFT corrections. It assumes that there are no invisible or exotic Higgs decays.
It is straightforward to use this method to compare the expectations for the precision in Higgs coupling measurement between linear and circular colliders. In Table 1, I present the expected 1 $\sigma$ precisions (in %) for LCF, FCC-ee, and CEPC I assume the same detection efficiencies and background subtractions for all colliders, based on the full simulation studies done for ILC. This makes clearer the differences in the results that are due to the different run plans, that is, from the integrated luminosities at the various energy stages and from beam polarization for LCF. For LCF, I show the results for the initial stage at 250 GeV with 3 ab-1 of data and for the full program of including 8 ab-1 of data at 550 GeV. For FCC-ee and CEPC, the main sources of data are the runs at 240 GeV, but the estimates given include also anticipated running at the top quark threshold and just above. For FCC-ee, I use the run plan in [16], with 10.8 ab-1 of data at 240 GeV. For CEPC, I use the run plan in [19], with 2 detectors and a total of 20 ab-1 of data at 240 GeV. All estimates include expected results from HL-LHC on the $H\to\gamma\gamma$ and $H\to\mu^{+}\mu^{-}$ decays. The final results for the expected precisions are remarkably similar. In all cases, the uncertainties are dominated by statistical errors, with systematic effects well under control.
For comparison, I present in Table 2 the corresponding results from the SMEFT fit carried out in the Briefing Book [11]. This fit follows a similar strategy with some differences in detail. More operators are included, but also more pieces of data constraining them. Most importantly, it assumes that the Higgs boson has no exotic decays. This table shows the same general trends and the essential equivalence of the full programs of the linear and circular Higgs factories for these coupling measurements.
[…]
A question often asked about Higgs coupling precision is: Yes, there are improvements over what is possible at LHC, but is it possible to access the effects of new particles? Given the success of the SM and the fact that, with a Higgs boson mass of 125 GeV, the SM could be valid up to the Planck scale, it is not possible to guarantee a discovery, either in direct particle searches or in precision studies. However, there are real oppportunites for discovery available. Some general classes of models giving sensitivity to new particles at the levels of precision given in Table 1 are reviewed in [28].
[…]
Figure 7: Relative deviations of the Higgs boson couplings in six diverse models of new physics. First row: 2-Higgs doublet models; Second row: Little Higgs models; Third row: a Composite Higgs model and a Scalar Singlet model. Error intervals shown are those for the ILC500. From [21]; see [26] for details on the specific models chosen.
Surveying models of BSM physics more broadly, more such opportunities are available. From model to model, the deviations occur in different Higgs boson coupling. Many different Higgs couplings will be measured with high precision, and so the pattern of deviations gives information about the higher energy theory that gave rise to them. Figure 7, from [21], illustrates this. The precision study of Higgs couplings gives us the opportunity not only to discover deviations from the SM but also to gain important clues as to its origin.
4 Study of the $Z$ boson at high statistics
The high luminosities available at lower energies at circular colliders make it possible to imagine a very high statistics program of measurements at the $Z$ pole, gathering about $5\times 10^{12}$ $Z$ decays (Tera-Z). Even with the lower luminosities available at linear colliders, it is possible to gather a sample of $5\times 10^{9}$ $Z$ decays, an improvement of more than 3 orders of magnitude over the LEP program, and with polarized beams (Giga-Z). It is interesting to discuss the potential for the discovery of BSM effects with these large data set.
One goal of this program repeats that of LEP: Test the SM by measuring $\sin^{2}\theta_{w}$ to high precision. In both programs, this can be measured from electron left-right asymmetry of the $Z$ coupling, $A_{e}$. In the Tera-Z program this is done using the production angle asymmetry of $Z\to\tau^{+}\tau^{-}$ decays; in the Giga-Z program, this is done using the asymmetry of the total hadronic cross sections using polarized beams. Both techniques reach uncertainties $\Delta\sin^{2}\theta_{w}/\sin^{2}\theta_{w}\sim 1.5\times 10^{-5}$, an improvement of more than an order of magnitude from the current uncertainty. In the Tera-Z program, one can also access $A_{\ell}$ from the forward-backward asymmetries in $\mu^{+}\mu^{-}$ and $b\overline{b}$ production. I will comment on these below.
Figure 8: A representative diagram that includes a dimension-6 SMEFT operator and modifies the $Z$ boson coupling to fermions at 1-loop order in electroweak interactions.
However, this program also contains a more powerful probe of Beyond-SM physics. The precisions are such that the measurements are sensitive to loop corrections from new heavy particles that are also subject to SMEFT suppressions $v^{2}/M^{2}\sim 1\%$[33, 34, 35]. A typical diagram including a dimension-6 SMEFT operator in a $Z$ radiative correction is shown in Fig. 8. For $M\sim 1$ TeV, this would be a deviation from the SM precision electroweak predictions of order $10^{-4}$. With the very high statistics available from a $Z$ pole program, these effects give sensitivity to new physics beyond the SM at the 10 TeV scale and beyond. The potential sensitivity is comparable to that available from measurements of heavy quark weak interactions. But these effects do not rely on flavor violation or violation of weak interaction universality. They are produced by generic modifications at the BSM energy scale.
Still, there are two qualifications that must be given to the promise of the previous paragraph. First, though new physics effects can come from a very wide range of operators, the strongest observable effects occur in a relatively small number of electroweak observables. The most important ones are: (1) the total width $\Gamma_{Z}$, (2) the hadron to lepton ratios $R_{e}$, $R_{\mu}$, $R_{\tau}$, (3) the ratio $R_{b}$ of $b$ decays to all hadronic decays, which specifically probes operators of the 3rd generation, and (4) $m_{W}$ or the $T$ parameter, which probe custodial symmetry violation and might also signal effects of the heavy top quark. In all cases, the effects would be tiny modifications of these quantities, which must be compared to high-precision predictions of the SM values.
This raises the second qualification: Can we produce the SM predictions with sufficient precision, and defend these in the case that an anomaly is observed? This is the question of the understanding and reliability of systematic uncertainties. The issues of raised by this question were clarified in the studies of Ayres Freitas for the 2026 ESPP update [37] and are discussed in Section 3.4 of the Briefing Book [11]. Freitas distinguished two sources of systematic error: first, the uncertainties from omitting terms beyond a given order in electroweak perturbation theory, second, the uncertainties in the modeling of physics events and the conversion of what is actually measured to ideal “pseudo-observables” to be compared to the results of calculation. I will give my own opinion on these sources, which differs from the attitude taken in the Briefing Book.
On the first issue, the truncation of perturbation theory, I have great confidence that theorists will achieve sufficiently accurate calculations. Essentially, precision electroweak observables and the electron structure functions must be computed to N3LO order with general nonzero masses. This is a major challenge. However, theorists have met a similarly difficult challenge in improving methods for perturbative QCD to provide N3LO and multijet cross sections needed for LHC. Also, this is a well-defined problem, so that, once it is solved with new insights, many groups will verify the results.
On the second issue, I have much less confidence. The sources of uncertainty here are come from non-perturbative effects as they are modeled in the widely used Monte Carlo programs. This is not a matter of lacking statistics; what we lack is an accurate theoretical understanding of the non-perturbative stage of hadronization. It is noteworthy, and somewhat depressing, that we still estimate systematic errors from hadronization from the differences between the results of PYTHIA and HERWIG. Today, there is no strategy improve this situation. We should treat uncertainties from this source more conservatively.
The authors of the Briefing Book estimated “conservative” and “aggressive” systematic uncertainties on the important precision electroweak measurements. The “conservative” estimates represent improvements, though not revolutionary ones, in the current state of the art. The “aggressive” uncertainties are, essentially, what is needed to overcome the barriers described in the previous paragraph. They include the improvement of uncertainties in nonperturbative hadronization by a factor 50 and setting to zero uncertainties that we currently do not know how to estimate. This seems to overlook the requirement of defending any claim of violation of the SM against the skepticism of our colleagues.
The “conservative” estimates of theory uncertainties are listed in Table 3, along with the statistical and experimental systematic uncertainties expected from the $Z$ pole programs of LCF and FCC-ee. It is important to note that these theory systematic uncertainties are large compared to the claimed experimental systematics, especially for the Tera-Z program. The resulting uncertainties on the $Z$ couplings to fermions are given in Table 4. In contrast, the the projections given in the Briefing Book for the uncertainties in Higgs couplings (Table 2), which are not at this very high level of precision, are quite robust to the choice of “conservative” or “aggressive” uncertainties. The differences are at the 20% level in the estimated uncertainty on the Higgs coupling for the $Z$ and $W$ and at the percent level for the other cases. This same robustness appears for the results on top quark interactions discussed in the following section.
observable Tera-Z Th. Systematics Giga-Z
$Z$ width:
   $\Gamma_{Z}$ 0.05 0.2 [2.2] 5
Ratios of BRs:
   $R_{e},R_{\mu}$, $R_{\tau}$ 0.3 1.2 [3.2] 13
   $R_{b},R_{c}$ 0.20/0.26 20/100 9/29
$\sin^{2}\theta_{w}$:
   $A_{e}(A_{LR})$ 19 19
   $A_{e}(\tau\ \mbox{pol})$ 14 27
   $A_{FB,b},A_{FB,c}$ 5/9 31/31
   $A_{FB,\mu},A_{FB,\tau}$ 20 230
$W$ mass:
   $m_{W}$ 0.3 0.7 3
Table 3: Expected systematic uncertainties for the major precision electroweak observables. Values for FCC-$e$ are taken from [36]; values from LCF are taken from [24]. Theory systematics are taken from [37]. All uncertainties are given as relative uncertainties in units of $10^{-5}$. Values in brackets are the uncertainties generated by the uncertainty in the measured value of $\alpha_{s}$, assumed to be $\Delta\alpha_{s}=10^{-4}$.
$Z$ coupling to: LCF 250 FCC-ee 240
$e_{L}$ 0.015 0.011
$e_{R}$ 0.021 0.015
$\tau_{L}$ 0.022 0.013
$\tau_{R}$ 0.030 0.019
$c_{L}$ 0.054 0.051
$c_{R}$ 0.082 0.071
$s_{L}$ 0.043 0.040
$s_{R}$ 0.54 0.34
$b_{L}$ 0.023 0.019
$b_{R}$ 0.26 0.33
Table 4: Expected uncertainties in the $Z$ couplings to fermions from the LCF 250 and FCC-ee 240 experimental programs from the SMEFT fit in the Briefing Book [11], under the “conservative” assumptions for theory systematic uncertainties. In both programs, the strongest constraints on these couplings come from the $Z$ pole measurements.
Even with the “conservative” theory systematic uncertainties, the power of the $Z$ pole measurements is very impressive. This is often presented as a table or graph of sensitivity scales $\Lambda$ for higher-dimension SMEFT operators. In (4), the coefficient of each dimension-6 operator is written as $c_{i}/\Lambda^{2}$, where $\Lambda^{2}$ is a mass scale. In an underlying BSM model, effects of new heavy particles generate corrections to the SM, and these are described by specific higher-dimension operators included in the SMEFT Lagrangian. In a fit to data with a Lagrangian including some number of these operators, the standard error $\sigma$ on each coefficient gives an idea of the masses of the new particles generating this operator to which that data set is sensitive. That scale would be roughtly $\Lambda\sim 1/\sqrt{\sigma}$ for operators generated by new strong interactions and $\Lambda/\sqrt{\alpha_{w}}\sim\sqrt{\sigma}$, given $\Lambda$ values a factor 6–10 smaller, for operators generated by electroweak radiative corrections.
This can be viewed in two ways. First, one can include all dimension-6 operators that respect basic symmetries of the theory. The fit in the Briefing Book assumes SM gauge symmetry, CP conservation, and a $U(2)$ flavor symmetry for the lighter two fermion generations. This leads to a Lagrangian with 124 dimension-6 operators. Of these, the 84 operators with unsuppressed couplings to precision observables at the tree or 1-loop level were included in the analysis. For each operator, the uncertainty in the operator coefficient given by the global if can be converted to a mass scale using the relations just described. This estimates the mass scale for which an observed deviation in the data can be shown to be induced by that operator. From the electroweak observables, this fit gives sensitivities to new particle masses in the few-TeV region. I will give some examples below in Table 5 in the following section.
However, one can also probe the data in another way, by including only one dimension-6 operator in the fit. This would represent a particular mode of violation of the SM. In that case, the uncertainty in the SMEFT coefficient for that operator and the associated $\Lambda$ scale estimates the sensitivity to new physics from all possible operators creating similar effects. That is, these single-operator fits estimate in more general way the mass scale for to which the data is sensitive. In that case, the fits to expected precision electroweak data give sensitivity to scales $\Lambda\sim 20-60$ TeV. Specific values for different top quark operators are also given in Table 5.
Deviations from the SM predictions for precision electroweak observables could well be the first evidence the that the SM must be corrected by new physics at a high mass scale. This is a capability within the reach of both circular and linear Higgs factories.
5 Higgs measurements above the top quark threshold
Finally, I will discuss the Higgs factory program at energies above the top quark threshold. This region is often omitted from discusssions of the physics of Higgs factories, but in fact it allows additional important measurements whose results are essential to the Higgs story.
First, running at high energy will allow us to redo the measurements of the Higgs couplings listed in Table 1. As I pointed out above, the $WW$ fusion reaction becomes the dominant source of Higgs boson prodiuction above 400 GeV, so the new Higgs sample will be dominated by this new reaction, with different experimental issues and sources of uncertainty. This will be an important check on any anomalies discovered in the program at the Higgs threshold. Assuming compatibility of the two sets of results, the improvement in the uncertainties on Higgs couplings has already been shown in Table 1.
Second, the top quark Yukawa coupling can be measured at CM energies of 550 GeV and above in the reaction $e^{+}e^{-}\to t\overline{t}H$. From ILC and CLIC studies, we expect an uncertainty of 2.8% with 8 ab-1 of data at 550 GeV and 1% at with 8 ab-1 of data at 1 TeV [24]. This latter value matches the statistical uncertainty on the top quark Yukawa coupling eventually expected from the FCC-hh  [16], but in an environment in which the systematic errors are subdominant.
Figure 9: Absolute precision on the deviation of the Higgs self-coupling from its SM value ($\Delta\kappa_{\lambda}$) expected from the measurement of Higgs pair production in the ZHH and $W$-fusion processes at 550 GeV, showing the complementary information from the two production modes, from [24].
Third, a Higgs factory operating at 550 GeV and above can measure Higgs pair production, which directly accesses the Higgs boson self-coupling. An $e^{+}e^{-}$ collider at 550 GeV can measure Higgs pair production in two reactions, $e^{+}e^{-}\to ZHH$ and the $WW$ fusion reaction $e^{+}e^{-}\to\nu\overline{\nu}HH$. With 8 ab-1, the expected precision in the measurement of the Higgs self-coupling is 11%. This is illustrated in Fig. 9 [24]. However, this is not the most important conclusion from this figure. Notice that the combined uncertainty in the deviation from the SM ($\Delta\kappa_{\lambda}$) is uniform as a function of the true underlying value of the self-coupling. The reason for this is that the amplitude for Higgs pair production is a sum of the contribution from the self-coupling and contributions due to other SM processes. In the $ZHH$ reaction, these amplitudes are in constructive interference; in the $WW$ fusion reaction, they are in destructive interference. For a positive deviation of $\kappa_{\lambda}$, we should observe an increased value of the $ZHH$ cross section and a decreased value of the $WW$ fusion cross section in the same experiment. This observation is possible only at an $e^{+}e^{-}$ collider. Observing both effects might be necessary to provide a credible discovery that the Higgs self-coupling differs from the SM prediction. The figure also shows the improvement in precision (to about 5% over most of the region) available with an additional 8 ab-1 at a CM energy of 1 TeV.
[…]
More generally, the CM energy dependence , angular dependence, and beam-polarization dependence of the reaction $e^{+}e^{-}\to t\overline{t}$ can all be sensitive to modifications induced by dimension-6 operators. With enough observables measured, it is possible to distinguish the effects of different operators. This has been tested in full-simulation studies carried out by the ILC and CLIC groups for event samples at 380 GeV, 550 GeV, and 1 TeV. These studies are described in [38, 39], with some updated results presented in [24]. The method used is to fit to simulation data the predictions from a SMEFT Lagrangian with 21 dimension-6 operators including the 11 most important operators bilinear in top quark fields. The covariance matrix obtained for the dimension-6 operator coefficients was then used as an input to the larger 84-operator global fit carried out for the Briefing Book. The resulting uncertainties for the 11 top quark operators are given in Table 5.
operator LCF 250 FCC-ee 240 FCC-$e$ LCF 550 LCF1000
electroweak couplings:
   $C_{\phi Q}^{-}$ 0.65 / 27.5 0.84 / 30.8 1.12 / 34.4 1.37/ 29.7 2.48 / 34.1
   $C_{\phi Q}^{3}$ 1.02 / 37.9 1.27 / 41.3 1.66 / 44.2 2.94 / 39.0 4.43 / 41.9
   $C_{\phi t}$ 0.86 / 18.5 0.95 / 22.4 1.46 / 27.8 1.44 / 23.5 2.49 / 30.2
Yukawa coupling:
   $C_{t\phi}$ 0.70 / 2.63 0.76 / 2.92 0.85 / 2.99 1.20 / 2.91 1.44 / 3.07
4-fermion operators:
   $C_{Q\ell}^{-}$ 0.98 / 30.5 1.21 / 37.8 4.01 / 45.8 7.49 / 65.2 15.5 / 95.2
   $C_{Q\ell}^{3}$ 1.36 / 42.9 1.67 / 50.6 3.41 / 62.3 8.93 / 85.1 18.1 / 127.
   $C_{te}$ 0.78 / 14.1 1.03 / 24.4 4.07 / 26.2 4.10 / 45.3 15.1 / 61.5
   $C_{t\ell}$ 1.28 / 14.3 1.38 / 24.8 5.29 / 27.7 5.90 / 42.6 20.3 / 56.6
   $C_{Qe}$ 2.88 / 32.3 3.00 / 28.8 3.63 / 31.4 10.9 / 62.4 18.3 / 85.0
magnetic moments:
   $C_{tW}$ 1.25 / 26.0 1.49 / 33.8 4.68 / 41.3 5.75 /40.8 6.67 /44.4
   $C_{tZ}$ 1.16 / 23.2 1.20 / 29.7 3.99 / 36.1 7.42 /37.6 8.69 / 40.7
[…]
This Table requires some explanation. The operators included in the left-hand column are: the 3 operators that shift the left- and right-handed top quark electroweak couplings, the operator that shifts the top quark Yukawa coupling, the 5 operators that provide 4-fermion $e^{+}e^{-}t\overline{t}$ contact interactions, and the 2 operators that shift the top quark $W$ and $Z$ magnetic moments. All of these operators can be induced by BSM top quark physics. I have discussed the shift in the electroweak couplings above. The 4-fermion operators can be induced by constituent exchange in composite top quark models, or by the Kaluza-Klein vector bosons in a higher-dimensional model of Higgs compositeness. Four-fermion $e^{+}e^{-}t\overline{t}$ operators can also be induced as NLO corrections to order-1 4-top contact interactions. It is important to recognize that the SM is violated by BSM effects, more important to demonstrate that this effect is specifically associated with the top quark, and still more valuable to distinguish between the electroweak couplings and 4-fermion effects, which are characteristic of different classes of models.
The values given in the table are the uncertainties in the operator coefficients interpreted, as I explained at the end of Section 4, as sensitivities to large values of the new physics scales $\Lambda$ in the SMEFT Lagrangian. The first number of each pair is the sensitivity of this particular operator from the 84-operator fit. The second number is the sensitivity from a fit that includes only this operator. This tests the ability of this operator perturbation to signal a deviation from the SM. To the extent that these numbers are of comparable size, the fit indicates that the data is able to pick out that particular operator as the one responsible for an observed deviation.
The single-operator sensitivities in the first two columns are already very impressive. These are mainly obtained from the $Z$ pole programs of the LCF and FCC-ee. As I noted in Section 4, the Tera-Z program gives higher sensitivities, but only to the extent that the uncertainties presented Table 4 are smaller. The next two columns include the first stages beyond the top quark threshold. The sensitivites from direct $t\overline{t}$ observations surpass those from the $Z$ pole measurements, and more substantially as the CM energy is raised. The specificity of the operator identifications is still low, because, with only one CM energy, there is a degeneracy in the effects from the top quark electroweak modifications and the top 4-fermion couplings. However, the electroweak modifications to the cross section stay at a constant size as a function of CM energy, while the effects of 4-fermion operators increase as $E_{CM}^{2}$. Then, adding a data set at 1 TeV clearly distinguishes the two effects, giving marked increases in the ability of the fit to pick out the specific top quark operator responsible. Also, to the extent that the new physics induces 4-fermion interactions, the 1 TeV data substantially improves the overall sensitivity to BSM physics.
At the CM energies where top quark observations dominate the SMEFT fit results, those results again become relatively insensitive to theory systematics, giving very similar results from the “conservative” and “aggressive” assumptions discussed in Section 4. I commented there that observations of tiny deviations in the precision electroweak observables may not have sufficient power to convince skeptics that the SM is in fact violated. For BSM physics associated with the third generation, observation of deviations from the SM predictions of the top quark interactions would strongly confirm any effect suggested by the precision electroweak program. This statement generalizes to other source of BSM physics, since high energy $e^{+}e^{-}$ experiments will look for deviations in the complete set of final state available through 4-fermion contact interactions.
This completes our survey of the physics opportunities of Higgs factories. These include the ability to discover effects from extended Higgs bosons, top quark partners, and other BSM particle well beyond the direct discovery limits of the HL-LHC, and the ability to probe for new BSM interactions at scales of 40 TeV and above. For lower energies, circular colliders have an advantage due to their higher luminosities, but this advantage is surprisingly modest. This is quantified above in Tables 14.
The measurements described in this section on the top quark Yukawa coupling, the Higgs self-coupling, and the search for compositeness in top quark interactions add important capabilities that are available at CM energies of 550 GeV and above, a region that can be probed only with linear collider Higgs factories. Probably the most important feature of measurements in this energy region is that they can provide reactions distinct from those at lower energy that are sensitive to the same models of BSM physics, giving the ability to make multiple independent measurements of suggested new physics effects. Without these, it will be difficult to fulfill the high burden of proof for a discovery based on precision measurement. The FCC-ee at CERN has powerful capabilities for discovery. But we may well need this higher-energy program also to build a truly persuasive case for new physics beyond the SM.
6 The challenge for now
The construction of a Higgs factory to begin operation just after the completion of the HL-LHC program is important for our understanding of physics, but it is even more important for the early career scientists in our community. The HL-LHC, planned to operate in 2030, offers exciting opportunites for the discovery of new particles, especially for color-neutral states such as extended Higgs states and electroweakinos with masses in the range of 1 TeV. But we can already see the limits of this program and the need to look beyond it. A new major collider takes 10 years to construct. This will be the dominant machine at the peak of your career. It is time to start thinking about it now.
This is going to be fun. The environment of an $e^{+}e^{-}$ collider is very different from that of the LHC. There is minimal radiation load on the detectors, no underlying event, no pileup. The $e^{+}e^{-}$ annhilation events are simpler than those at the LHC. Not only are the particle multiplicities lower, but there is no forward fragmentation lost down the beam pipe. This is an environment in which it is possible to make vertex, tracking, and calorimetry measurements that are close to perfect. But – this point is very important – this does not make experimentation easier, because great improvements in event reconstruction are need to satisfy the goals of high precision measurements.
It is likely that the Higgs factories will provide the only chance that you will have in career to design a new collider detector from the ground up. This detector should incorporate new technologies and should be adapted for heavy use of Machine Learning and AI. This is a fascinating challenge. It is also one for which you are much better prepared than your professors.
As a first step, you should study the work that has already been done to design detectors specifically adapted to make high precision measurements at $e^{+}e^{-}$ colliders. I recommend, in particular, close study of the ILD Interim Design Report [40], which gives a thoughtful approach to precision measurement with currently available technologies. Detectors meant to operate at Higgs factories in the 2040’s should go even further to incorporate new ideas. A review of such new technologies can be found in the 2023 community report [41], though already this document is somewhat out of date.
Let me highlight as examples three directions that are promising areas for detector R&D. The first of these is in tracking, with the use of Monolithic Active Pixel Sensors (MAPS). These detectors incorporate both sensors and readout electronics in wafers with minimal material to decrease contributions to track pointing resolution from multiple scattering. The pixel size is less than $25~\mu$; there is no need for bump bonding. The wafers can be thinned to less than $50~\mu$, at which point they are bendable and require only a simple carbon-fiber support structure. In principle, such detectors can incorporate timing and, at the same time, operate at very low power, avoiding the material of a cryostatThis goal may be challenging to achieve with the high event rates of the Tera-Z program.. MAPS are already incorporated in the design of the ALICE vertex detector upgrade for HL-LHC [42], though this design must still deal with a high-radiation environment. The possible adaptation of this technology to $e^{+}e^{-}$ experimentation is still begin explored.
The second is in calorimetry. From the first studies for calorimeters for linear $e^{+}e^{-}$ collisions, it was envisioned that, in the $e^{+}e^{-}$ environment, one could use particle flow calorimetry together with a high granularity calorimeter to improve calorimeter resolution to $25\%/\sqrt{E}$ and below. The calorimeter would be based on a tungsten absorber with small Molière radius interleaved with silicon sensors [43]. This technology has been investigated by the CALICE collaboration; the current status is described in  [44]. This strategy is incorporated in the CMS High-Granularity forward calorimeter for the HL-LHC upgrade [45].
An alternative technology would be dual readout calorimetry, in which one would collect both scintillation and Cherenkov light signals from calorimetry modules, allowing separation of the charged and neutral components of a hadronic shower [46]. This method is now undergoing rapid development, with the inclusion of longitudinal segmentation, timing, and particle flow analysis strategies [47, 48]. Up to now, incorporation of Machine Learning in either technique has not led to large improvements in resolution, but it is surely just a matter of time before those ideas have their effect.
The third is in particle ID. Recently, there have been major advances in particle ID in the LHC environment from the use of machine learning. A very influential step was made in the ParticleNet architecture of Qu and Gouskos [49]. An application of this approach to full-simulation Higgs factory data by the CEPC group is reported in [50, 51]. These studies are very promising for the ability to measure $H\to s\overline{s}$ at 250 GeV. Here also, the level of improvement possible in the $e^{+}e^{-}$ environment is still being explored.
This brings up a more general question that I feel is very important. So far, all of the AI applications to event analysis have been done in the context of existing designs and hardware, both for LHC and for Higgs factories. But to develop a develop a future detector, as AI gives us more capabilities, the hardware design should co-evolve with the analysis software to use the new methods most effectively. Detectors for the 2040’s will have some level of intelligence on every sensor. Does this push us to higher granularity to gain more information or to lower granularity, with local analysis of the waveforms produced by the sensors? How will we read out highly intelligent sensors? How will we incorporate timing information? How to these solutions interact with more traditional restrictions to minimize material and heating? The answers to these questions may well lead us to radically new ideas about detector design and performance.
This is a fascinating challenge for early-career experimental physicists. The people with the best ideas will be the leaders of collider physics in the Higgs factory era. These questions need your attention now. It is also time to demand serious R&D funding to build prototypes that will validate or challenge your most innovative concepts.
7 Conclusions
In this lecture, I have discussed the following points:
  • The central role of the Higgs boson our understanding of particle physics, and the importance of studying the Higgs boson at higher levels of precision
  • The opportunities for the discovery of signals of physics beyond the Standard Model in precision studies of Higgs couplings and higher levels of precision in electroweak measurements.
  • The importance of precision measurements of properties of the Higgs boson and the top quark that require $e^{+}e^{-}$ CM energies well above the $t\overline{t}$ threshold.
  • The challenge of constructing Higgs factory detectors of ultimate precision to carry out these measurements.
For early-career physicists, the Higgs factory era will be a time of adventure and discovery. Our community needs to bring this era to reality on as fast a timeline as possible.
ACKNOWLEDGEMENTS
I am grateful to many colleagues for discussions of the physics capabilities of $e^{+}e^{-}$ Higgs factories over the course of the European Strategy study. I thank especially Alain Blondel, Jorge de Blas, Ayres Freitas, Jenny List, Kamal Maayergi, Matthew McCullough, Patrick Meade, Roman Pöschl, Aidan Robson, Ariel Schwartzman, Ben Stefanek, Junping Tian, Caterina Vernieri, and Devin Walker. This work was supported by the US Department of Energy under contract DE–AC02–76SF00515.

### v3

@@
 We need to learn more about this particle, and we have the resources to gain that knowledge.
-We will learn much from the High Luminosity phase of the LHC, but, beyond this, we will be able to measure the properties of the Higgs boson with high precision from a future “Higgs Factory”, an $e^{+}e^{-}$ collilder operating at energies from the $Z$ boson resonance to the top quark threshold and above.
+We will learn much from the High Luminosity phase of the LHC, but, beyond this, we will be able to measure the properties of the Higgs boson with high precision from a future “Higgs Factory”, an $e^{+}e^{-}$ collider operating at energies from the $Z$ boson resonance to the top quark threshold and above.
 This has been recognized in the latest strategy statements from all of the major world regions in particle physics.
@@
 From the 2023 US High Energy Physics Advisory Panel P5 report: “The next step is to use electron and positron beams to construct a Higgs factory, which would allow precision measurements of the Higgs boson properties and searches for exotic decays, possibly into dark matter” [8].
-In China, the Circular Electron-Positron Collilder (CEPC) was named the highest priority next project in particle and nuclear physics in the first round of CAS deliberations [9], though eventually the CAS chose to endorse a less ambitious collider.
+In China, the Circular Electron-Positron Collider (CEPC) was named the highest priority next project in particle and nuclear physics in the first round of CAS deliberations [9], though eventually the CAS chose to endorse a less ambitious collider.
 Very recently, the European Strategy Group has recommended the FCC-ee, a circular Higgs factory, as the next flagship CERN project [10].
@@
 Section 2 reviews the central role of the Higgs boson that motivates these measurements.
-In Sections 3–5, I discuss Higgs measurements in three different programs avaiable at Higgs factories.
+In Sections 3–5, I discuss Higgs measurements in three different programs available at Higgs factories.
 Section 3 discusses the program at the peak of the $e^{+}e^{-}\to HZ$ cross section at 240-250 GeV in the $e^{+}e^{-}$ center of mass.
@@
 1, from CMS, and a similar figure from ATLAS, both published in Nature [3, 4].
-This data demonstrates that the Higgs boson vacuum expectation value is the main source of mass for the particles shown.
+These data demonstrate that the Higgs boson vacuum expectation value is the main source of mass for the particles shown.
 Many people interpret these beautiful results as proof that the observed Higgs boson precisely follows the Standard Model.
@@
 We will see how these differences play out as we consider the various energy stages of $e^{+}e^{-}$ Higgs factory operation.
-Figure 3: Cross sections for Higgs boson production in $e^{+}e^{-}$ annhilation reactions as a function of CM energy, showing also the dominant Feynman diagrams for the reactions $e^{+}e^{-}\to ZH$ and $e^{+}e^{-}\to\nu\overline{\nu}H$.
+Figure 3: Cross sections for Higgs boson production in $e^{+}e^{-}$ annihilation reactions as a function of CM energy, showing also the dominant Feynman diagrams for the reactions $e^{+}e^{-}\to ZH$ and $e^{+}e^{-}\to\nu\overline{\nu}H$.
 Figure 4: Determination of the Higgs boson mass at 250 GeV from measurement of the $Z$ recoil mass in $e^{+}e^{-}\to ZH$ with $Z\to\mu^{+}\mu^{-}$, from [25].
@@
 The dominant cross sections for Higgs boson production at $e^{+}e^{-}$ colliders are Higgstrahlung, $e^{+}e^{-}\to ZH$, and $WW$ fusion, $e^{+}e^{-}\to\nu\overline{\nu}H$.
-The cross sections for these reactions in the CM energy range of Higgs factories is shown in Fig.
+The cross sections for these reactions in the CM energy range of Higgs factories are shown in Fig.
 3.
@@
 It is straightforward to use this method to compare the expectations for the precision in Higgs coupling measurement between linear and circular colliders.
-In Table 1, I present the expected 1 $\sigma$ precisions (in %) for LCF, FCC-ee, and CEPC I assume the same detection efficiencies and background subtractions for all colliders, based on the full simulation studies done for ILC.
-This makes clearer the differences in the results that are due to the different run plans, that is, from the integrated luminosities at the various energy stages and from beam polarization for LCF.
+In Table 1, I present the expected 1 $\sigma$ precisions (in %) for LCF, FCC-ee, and CEPC.
+I assume the same detection efficiencies and background subtractions for all colliders, based on the full simulation studies done for ILC.
+This makes clearer that the differences in the results that are due to the different run plans, that is, from the integrated luminosities at the various energy stages and from beam polarization for LCF.
 For LCF, I show the results for the initial stage at 250 GeV with 3 ab-1 of data and for the full program of including 8 ab-1 of data at 550 GeV.
@@
 Given the success of the SM and the fact that, with a Higgs boson mass of 125 GeV, the SM could be valid up to the Planck scale, it is not possible to guarantee a discovery, either in direct particle searches or in precision studies.
-However, there are real oppportunites for discovery available.
+However, there are real opportunities for discovery available.
 Some general classes of models giving sensitivity to new particles at the levels of precision given in Table 1 are reviewed in [28].
@@
 On the second issue, I have much less confidence.
-The sources of uncertainty here are come from non-perturbative effects as they are modeled in the widely used Monte Carlo programs.
+The sources of uncertainty here come from non-perturbative effects as they are modeled in the widely used Monte Carlo programs.
 This is not a matter of lacking statistics; what we lack is an accurate theoretical understanding of the non-perturbative stage of hadronization.
@@
 The resulting uncertainties on the $Z$ couplings to fermions are given in Table 4.
-In contrast, the the projections given in the Briefing Book for the uncertainties in Higgs couplings (Table 2), which are not at this very high level of precision, are quite robust to the choice of “conservative” or “aggressive” uncertainties.
+In contrast, the projections given in the Briefing Book for the uncertainties in Higgs couplings (Table 2), which are not at this very high level of precision, are quite robust to the choice of “conservative” or “aggressive” uncertainties.
 The differences are at the 20% level in the estimated uncertainty on the Higgs coupling for the $Z$ and $W$ and at the percent level for the other cases.
@@
 In a fit to data with a Lagrangian including some number of these operators, the standard error $\sigma$ on each coefficient gives an idea of the masses of the new particles generating this operator to which that data set is sensitive.
-That scale would be roughtly $\Lambda\sim 1/\sqrt{\sigma}$ for operators generated by new strong interactions and $\Lambda/\sqrt{\alpha_{w}}\sim\sqrt{\sigma}$, given $\Lambda$ values a factor 6–10 smaller, for operators generated by electroweak radiative corrections.
+That scale would be roughly $\Lambda\sim 1/\sqrt{\sigma}$ for operators generated by new strong interactions and $\Lambda/\sqrt{\alpha_{w}}\sim\sqrt{\sigma}$, for operators generated by electroweak radiative corrections, giving $\Lambda$ values a factor 6–10 smaller.
 This can be viewed in two ways.
@@
 Finally, I will discuss the Higgs factory program at energies above the top quark threshold.
-This region is often omitted from discusssions of the physics of Higgs factories, but in fact it allows additional important measurements whose results are essential to the Higgs story.
+This region is often omitted from discussions of the physics of Higgs factories, but in fact it allows additional important measurements whose results are essential to the Higgs story.
 First, running at high energy will allow us to redo the measurements of the Higgs couplings listed in Table 1.
-As I pointed out above, the $WW$ fusion reaction becomes the dominant source of Higgs boson prodiuction above 400 GeV, so the new Higgs sample will be dominated by this new reaction, with different experimental issues and sources of uncertainty.
+As I pointed out above, the $WW$ fusion reaction becomes the dominant source of Higgs boson production above 400 GeV, so the new Higgs sample will be dominated by this new reaction, with different experimental issues and sources of uncertainty.
 This will be an important check on any anomalies discovered in the program at the Higgs threshold.
@@
 Four-fermion $e^{+}e^{-}t\overline{t}$ operators can also be induced as NLO corrections to order-1 4-top contact interactions.
-It is important to recognize that the SM is violated by BSM effects, more important to demonstrate that this effect is specifically associated with the top quark, and still more valuable to distinguish between the electroweak couplings and 4-fermion effects, which are characteristic of different classes of models.
 The values given in the table are the uncertainties in the operator coefficients interpreted, as I explained at the end of Section 4, as sensitivities to large values of the new physics scales $\Lambda$ in the SMEFT Lagrangian.
@@
 The next two columns include the first stages beyond the top quark threshold.
-The sensitivites from direct $t\overline{t}$ observations surpass those from the $Z$ pole measurements, and more substantially as the CM energy is raised.
+The sensitivities from direct $t\overline{t}$ observations surpass those from the $Z$ pole measurements, and more substantially as the CM energy is raised.
 The specificity of the operator identifications is still low, because, with only one CM energy, there is a degeneracy in the effects from the top quark electroweak modifications and the top 4-fermion couplings.
@@
 Also, to the extent that the new physics induces 4-fermion interactions, the 1 TeV data substantially improves the overall sensitivity to BSM physics.
+This study addresses a number of goals.
+It is important to recognize that the SM is violated by BSM effects, more important to demonstrate that this effect is specifically associated with the top quark, and still more valuable to distinguish between the electroweak couplings and 4-fermion effects, which are characteristic of different classes of models.
+Table 5 shows that the program of high-energy running of an $e^{+}e^{-}$ collider will address all of these questions.
 At the CM energies where top quark observations dominate the SMEFT fit results, those results again become relatively insensitive to theory systematics, giving very similar results from the “conservative” and “aggressive” assumptions discussed in Section 4.
@@
 For BSM physics associated with the third generation, observation of deviations from the SM predictions of the top quark interactions would strongly confirm any effect suggested by the precision electroweak program.
-This statement generalizes to other source of BSM physics, since high energy $e^{+}e^{-}$ experiments will look for deviations in the complete set of final state available through 4-fermion contact interactions.
+This statement generalizes to other source of BSM physics, since high energy $e^{+}e^{-}$ experiments will look for deviations in the complete set of final states available through 4-fermion contact interactions.
 This completes our survey of the physics opportunities of Higgs factories.
@@
 The construction of a Higgs factory to begin operation just after the completion of the HL-LHC program is important for our understanding of physics, but it is even more important for the early career scientists in our community.
-The HL-LHC, planned to operate in 2030, offers exciting opportunites for the discovery of new particles, especially for color-neutral states such as extended Higgs states and electroweakinos with masses in the range of 1 TeV.
+The HL-LHC, planned to operate in 2030, offers exciting opportunities for the discovery of new particles, especially for color-neutral states such as extended Higgs states and electroweakinos with masses in the range of 1 TeV.
 But we can already see the limits of this program and the need to look beyond it.
@@
 There is minimal radiation load on the detectors, no underlying event, no pileup.
-The $e^{+}e^{-}$ annhilation events are simpler than those at the LHC.
+The $e^{+}e^{-}$ annihilation events are simpler than those at the LHC.
 Not only are the particle multiplicities lower, but there is no forward fragmentation lost down the beam pipe.
@@
 This strategy is incorporated in the CMS High-Granularity forward calorimeter for the HL-LHC upgrade [45].
-An alternative technology would be dual readout calorimetry, in which one would collect both scintillation and Cherenkov light signals from calorimetry modules, allowing separation of the charged and neutral components of a hadronic shower [46].
+An alternative technology would be dual readout calorimetry, in which one would collect both scintillation and Cherenkov light signals from calorimetry modules, allowing separation of the electromagnetic and non-electromagnetic components of a hadronic shower [46].
 This method is now undergoing rapid development, with the inclusion of longitudinal segmentation, timing, and particle flow analysis strategies [47, 48].
@@
 How will we incorporate timing information?
-How to these solutions interact with more traditional restrictions to minimize material and heating?
+How do these solutions interact with more traditional restrictions to minimize material and heating?
 The answers to these questions may well lead us to radically new ideas about detector design and performance.
@@
 I am grateful to many colleagues for discussions of the physics capabilities of $e^{+}e^{-}$ Higgs factories over the course of the European Strategy study.
-I thank especially Alain Blondel, Jorge de Blas, Ayres Freitas, Jenny List, Kamal Maayergi, Matthew McCullough, Patrick Meade, Roman Pöschl, Aidan Robson, Ariel Schwartzman, Ben Stefanek, Junping Tian, Caterina Vernieri, and Devin Walker.
+I thank especially Alain Blondel, Jorge de Blas, Ayres Freitas, Jenny List, Kamal Maayergi, Matthew McCullough, Patrick Meade, Roman Pöschl, Aidan Robson, Ariel Schwartzman, Ben Stefanek, Junping Tian, Caterina Vernieri, Devin Walker, and Filip Żarnecki.
 This work was supported by the US Department of Energy under contract DE–AC02–76SF00515.